OpenStudy (anonymous):

Solve for x in the proportion 4 times x squared over the quantity 8 times x squared minus 8 times x equals 2 over x. x = 2 and x = −2 x = 2 x = 16 and x = −16 x = −16

OpenStudy (anonymous):

@xoya

OpenStudy (anonymous):

It's x=2 :P

OpenStudy (anonymous):

\[\frac{ 4x^2 }{ 8x^2 - 8x }=\frac{ 2 }{ x }\] your going to want to multiply away the fractions, So multiply both sides by x, and both by 8x^2 - 8x, Can you work out the rest from there?

OpenStudy (anonymous):

or xoya can give you the answer...